# Artificial Intelligence for Microbiology and Microbiome Research

### v1

@@ 3.8.4 Phage therapy
As the most abundant organisms in the biosphere, bacteriophages (a.k.a. phages) are viruses that specifically target bacteria and archaea. They play a significant role in microbial ecology by influencing bacterial populations, gene transfer, and nutrient cycles. Moreover, they can be an alternative to antibiotics and hold the potential therapeutic ability for bacterial infections [343, 344, 345, 346].
@@
Phage identification. Many computational tools have been developed to identify bacteriophage sequences in metagenomic datasets [347]. They can be roughly grouped into two classes: (1) alignment-based (or database-based) methods, e.g., MetaPhinder [348], VIBRANT [349], and VirSorter2 [350]; (2) alignment-free (or learning-based) methods, e.g., VirFinder [351], PPR-meta [154], Seeker [352], DeepVirFinder [353], and PhaMer [354]. Alignment-based methods typically use a large number of sequences of references and utilize DNA or protein sequence similarity as the main feature to distinguish phages from other sequences. Their limitations are evident. Firstly, bacterial contigs may align with multiple phage genomes, potentially resulting in false-positive phage predictions. Secondly, novel or highly diverged phages may not have significant alignments with the selected phage protein families, which can lead to low sensitivity in identifying new phages. Alignment-free methods can overcome those limitations via machine learning or deep learning techniques. Those methods learn the features of the sequence data and are mainly classification models with training data consisting of both phages and bacteria. Some classification models use manually extracted sequence features such as k-mers, while others use deep learning techniques to automatically learn features. For example, VirFinder uses $k$-mers to train a logistic regression model for phage identification. Seeker (or DeepVirFinder) uses one-hot encoding to represent the sequence data and trains an LSTM (or CNN) to identify phages, respectively. PhaMer leverages the start-of-the-art language model, the Transformer, to conduct contextual embedding for phage contigs. It feeds both the protein composition and protein positions from each contig into the Transformer, which learns the protein organization and associations to predict the label for test contigs. It has been shown that PhaMer outperforms VirSorter, Seeker, VirFinder, DeepVirfinder, and PPR-meta.
@@
Recently, a hybrid method called INHERIT was developed. INHERIT (IdentificatioN of bacteriopHagEs using deep RepresentatIon model with pre-Training) naturally ‘inherits’ the characteristics from both alignment-based and alignment-free methods [355]. In particular, INHERIT uses pre-training as an alternative way of acquiring knowledge representations from existing databases, and then uses a BERT-style deep learning framework to retain the advantage of alignment-free methods. The independent pre-training strategy can effectively deal with the data imbalance issue of bacteria and phages, helping the deep learning framework make more accurate predictions for both bacteria and phages. The deep learning framework in INHERIT is based on a novel DNA sequence language model: DNABERT [60], a pre-trained bidirectional encoder representation model, which can capture global and transferrable understanding of genomic DNA sequences based on up and downstream nucleotide contexts. It has been demonstrated that INHERIT outperforms four existing state-of-the-art approaches: VIBRANT, VirSorter2, Seeker, and DeepVirFinder. It would be interesting to compare the performance of INHERIT and PhaMer.
@@
Phage-host interaction prediction. Phages can specifically recognize and kill bacteria, which leads to important applications in many fields. Screening suitable therapeutic phages that are capable of infecting pathogens from massive databases has been a principal step in phage therapy design. Experimental methods to identify phage-host interactions (PHIs) are time-consuming and expensive; using high-throughput computational methods to predict PHIs is therefore a potential substitute. There are two types of computational methods for PHI prediction. One is alignment-based. We explicitly align the viral and bacterial whole-genome sequences and acquire matched sequences to indicate PHI. The other is alignment-free. We compare nucleotide features and/or protein features extracted from viral and bacterial genomes, and predict PHI using machine learning. Each type of method has its pros and cons. A benchmark study ([361]) of those alignment-free machine learning methods demonstrated that GSPHI [362] and PHIAF [363] are the two best deep learning-based methods for PHI prediction. PHIAF is a deep learning method based on date augmentation, feature fusion, and the attention mechanism. It first applies a GAN-based data augmentation module, which generates pseudo-PHIs to alleviate the data scarcity issue. Then it fuses the features originating from DNA and protein sequences for better performance. Finally, it incorporates an attention mechanism into CNN to consider different contributions of DNA/protein sequence-derived features, which provides interpretability of the predictions. GSPHI is a novel deep learning method for PHI prediction with complementing multiple information. It first initializes the node representations of phages and target bacterial hosts via a word embedding algorithm (word2vec). Then it uses a graph embedding algorithm (structural deep network embedding: SDNE) to extract local and global information from the interaction network. Finally, it uses a multi-layer perceptron (MLP) with two hidden layers to detect PHIs.
@@
Very recently, AI techniques have been applied to discover novel phage lysins [378, 379]. DeepLysin is a unified software package to employ AI for mining the vast genome reservoirs for novel antibacterial phage lysins [378]. DeepLysin consists of two modules: the lysin mining module and the antibacterial activity prediction module. The input of the lysin mining module is assembled contigs. This module utilizes traditional blastP/protein sequence alignment-based methods to identify putative lysins. The second module estimates the antibacterial activity of the putative lysins identified by the first module. This module utilizes multiple AI techniques, such as Word2vec and an ensemble classifier that integrates five common classifiers to differentiate diverse and complex protein features. It ultimately applies Logistic Regression as a non-linear activation function to produce final activity predictions as scores ranging from 0 to 1, with higher scores indicating increased antibacterial activity. One limitation of DeepLysin is that four types of manually selected features (i.e., composition-based feature, binary profile-based feature, position-based feature, physiochemical based feature) need to be provided to the classifier. The feature selection procedure apparently heavily relies on domain knowledge.
@@ 4.1 Tradeoff between interpretability and complexity
Machine learning models, especially deep learning models, often suffer from high complexity and low interpretability, hindering their application in clinical decision-making. In addition, deep learning models typically have more than thousands of neural weights whose training requires large sample sizes and high computational resources. We anticipate that those deep learning models can reach better performance than traditional machine learning models as long as the sample size is enough. However, in most clinic-related studies, traditional models (e.g., Random Forest) are still widely used due to their ease of implementation, smaller sample size requirement, and better interpretability.
@@
To address the interpretability issue, two different approaches can be employed. One approach is to employ methods such as SHAP (SHapley Additive exPlanations) [393], LIME (Local Interpretable Model-agnostic Explanations) [394] to enhance the interpretability of black-box models. SHAP is a game-theoretic method used to explain the output of any machine learning model. It links optimal credit allocation to local explanations by leveraging Shapley values from game theory and their related extensions. LIME is a technique that approximates any black box machine learning model with a local, interpretable model to explain each individual prediction. By applying SHAP and LIME, we can gain insights into complex deep learning models, identify biases, and improve transparency, crucial for applications in microbiome research.
@@
The other approach is to employ “white-box” models. For instance, ReduNet [395] is a white-box deep network based on the principle of maximizing rate reduction. The authors argued that, at least in classification tasks, a key objective for a deep network is to learn a low-dimensional, linearly discriminative representation of the data. The effectiveness of this representation can be assessed by a principled measure from (lossy) data compression, i.e., rate reduction. Appropriately structured deep networks can then be naturally interpreted as optimization schemes designed to maximize this measure. The resulting multi-layer deep network shares key characteristics with modern deep learning architectures, but each component of ReduNet has a well-defined optimization, statistical, and geometric interpretation. Applying ReduNet to microbiome data would be an interesting attempt. Unlike ReduNet, MDITRE is a supervised deep learning method specifically designed for microbiome research. It takes a phylogenetic tree, microbiome time-series data, and host status labels to learn human-interpretable rules for predicting host status [396]. The model consists of five hidden layers that can be directly interpreted in terms of if-then rule statements. The first layer focuses on phylogenetic relationships by selecting taxa relevant to predicting host status. The second layer focuses on time by identifying relevant time windows for prediction. The following layers determine whether the data from selected taxa and time windows exceed specific learned thresholds, and subsequently combine these conditions to generate the final rules for prediction.
@@ 4.3 Benchmarking evaluations
As we mentioned in previous sections several times, benchmarking evaluations are typically lacking in microbiology and microbiome research. Currently, there is no standardized pipeline for benchmarking machine learning or deep learning methods in microbiology and microbiome research. To ensure reproducibility across studies, it’s critical to standardize data preprocessing, which includes consistent methods for data collection, bioinformatics pipelines, and the profiling of microbiome taxonomies. Additionally, if feature dimension reduction is needed, it must be unbiased, using standardized methods for feature selection or reduction that apply uniformly across studies. Importantly, feature engineering should only be applied to training data and later evaluated on test data to avoid data leakage or overfitting. Furthermore, the creation of publicly available, well-annotated benchmarking datasets (analogous to MNIST or ImageNet in computer science) would provide the microbiome research community with reliable tools to assess and compare different machine learning models. Such datasets would accelerate progress and provide a framework for objective evaluation of new computational methods. Some attempts have been made in this regard. For example, MicrobiomeHD is a standardized database that compiles human gut microbiome studies related to health and disease [404]. It contains publicly available 16S data from published case-control studies, along with associated patient metadata. The raw sequencing data for each study was obtained and processed using a standardized pipeline. The curatedMetagenomicData package is another excellent example of benchmark microbiome datasets. It offers uniformly processed human microbiome data, including bacterial, fungal, archaeal, and viral taxonomic abundances, as well as quantitative metabolic functional profiles and standardized participant metadata [405]. This comprehensive, curated collection of metagenomic data is well-documented and easily accessible, making it suitable for benchmarking machine learning methods.
@@
Establishing benchmark datasets is critical for advancing AI application in microbiology and microbiome research. Such datasets enable consistent, unbiased comparisons of algorithms and promote the development of robust predictive models. By providing standardized data, the research community can evaluate AI methods on a level playing field, ensuring reproducibility and transparency. Similar to the successful DREAM challenges in genomics, a community-driven effort to create public benchmarking datasets will foster collaboration, accelerate discovery, and establish best practices for AI approaches in microbiology and microbiome research. Collaborative input is vital for making this a reality.

### v2

@@ 3.8.5 Phage therapy
As the most abundant organisms in the biosphere, bacteriophages (a.k.a. phages) are viruses that specifically target bacteria. They play a significant role in microbial ecology by influencing bacterial populations, gene transfer, and nutrient cycles. Moreover, they can be an alternative to antibiotics and hold the potential therapeutic ability for bacterial infections [466, 467, 468, 469, 470].
@@
Phage identification. Many computational tools have been developed to identify bacteriophage sequences in metagenomic datasets [471]. They can be roughly grouped into two classes: (1) alignment-based (or database-based) methods, e.g., MetaPhinder [472], VIBRANT [473], and VirSorter2 [474]; (2) alignment-free methods, e.g., VirFinder [475], PPR-meta [139], Seeker [476], DeepVirFinder [477], and PhaMer [478]. Alignment-based methods typically use a large number of sequences of references and utilize DNA or protein sequence similarity as the main feature to distinguish phages from other sequences. Their limitations are evident. Firstly, bacterial contigs may align with multiple phage genomes, potentially resulting in false-positive phage predictions. Secondly, novel or highly diverged phages may not have significant alignments with the selected phage protein families, which can lead to low sensitivity in identifying new phages. Alignment-free methods can overcome those limitations via machine learning or deep learning techniques. Those methods learn the features of the sequence data and are mainly classification models with training data consisting of both phages and bacteria. Some classification models use manually extracted sequence features such as k-mers, while others use deep learning techniques to automatically learn features. For example, VirFinder uses $k$-mers to train a logistic regression model for phage identification. Seeker (or DeepVirFinder) uses one-hot encoding to represent the sequence data and trains an LSTM (or CNN) to identify phages, respectively. PhaMer leverages the start-of-the-art language model, the Transformer, to conduct contextual embedding for phage contigs. It feeds both the protein composition and protein positions from each contig into the Transformer, which learns the protein organization and associations to predict the label for test contigs. It has been shown that PhaMer outperforms VirSorter, Seeker, VirFinder, DeepVirfinder, and PPR-meta. geNomad is a hybrid framework that combines the strengths of alignment-free and alignment-based models for concurrent identification and annotation of both plasmids and viruses in sequencing data [479]. To achieve that, geNomad processes user-provided nucleotide sequences via two distinct branches. In the sequence branch (“alignment-free”), the inputs are one-hot encoded and passed through an IGLOO neural network, which evaluates them by identifying non-local sequence motifs. In the marker branch (“alignment-based”), the proteins encoded by the input sequences are annotated with markers specific to chromosomes, plasmids, or viruses. Here, the key idea behind the IGLOO neural network is to leverage the relationships between “non-local patches” sliced from feature maps generated by successive convolutions to effectively represent long sequences, allowing it to handle both short and long sequences efficiently, unlike traditional RNNs which struggle with very long sequences [480].
@@
Recently, a hybrid method called INHERIT was developed. INHERIT (IdentificatioN of bacteriopHagEs using deep RepresentatIon model with pre-Training) naturally ‘inherits’ the characteristics from both alignment-based and alignment-free methods [209]. In particular, INHERIT uses pre-training as an alternative way of acquiring knowledge representations from existing databases, and then uses a BERT-style deep learning framework to retain the advantage of alignment-free methods. The independent pre-training strategy can effectively deal with the data imbalance issue of bacteria and phages, helping the deep learning framework make more accurate predictions for both bacteria and phages. The deep learning framework in INHERIT is based on a novel DNA sequence language model: DNABERT [58], a pre-trained bidirectional encoder representation model, which can capture global and transferrable understanding of genomic DNA sequences based on up and downstream nucleotide contexts. It has been demonstrated that INHERIT outperforms four existing state-of-the-art approaches: VIBRANT, VirSorter2, Seeker, and DeepVirFinder. It would be interesting to compare the performance of INHERIT and PhaMer. Also, since DNABERT is per-trained on human genomes, it would be interesting to see if INHERIT can be further improved when using a BERT model pre-trained on prokaryotic genomes.
@@
Phage-host interaction prediction. Phages can specifically recognize and kill bacteria, which leads to important applications in many fields. Screening suitable therapeutic phages that are capable of infecting pathogens from massive databases has been a principal step in phage therapy design. Experimental methods to identify phage-host interactions (PHIs) are time-consuming and expensive; using high-throughput computational methods to predict PHIs is therefore a potential substitute. There are two types of computational methods for PHI prediction. One is alignment-based. We explicitly align the viral and bacterial whole-genome sequences and acquire matched sequences to indicate PHI. The other is alignment-free. We compare nucleotide features and/or protein features extracted from viral and bacterial genomes, and predict PHI using machine learning [487]. Each type of method has its pros and cons. For example, PHIAF [214] is a deep learning method based on date augmentation, feature fusion, and the attention mechanism. It first applies a GAN-based data augmentation module, which generates pseudo-PHIs to alleviate the data scarcity issue. Then it fuses the features originating from DNA and protein sequences for better performance. Finally, it incorporates an attention mechanism into CNN to consider different contributions of DNA/protein sequence-derived features, which provides interpretability of the predictions. GSPHI [488] is a deep learning method for PHI prediction with complementing multiple information. It first initializes the node representations of phages and target bacterial hosts via a word embedding algorithm (word2vec). Then it uses a graph embedding algorithm (structural deep network embedding: SDNE) to extract local and global information from the interaction network. Finally, it uses an MLP with two hidden layers to detect PHIs. PHPGAT [215] is a novel phage–host prediction model. It integrates a multimodal heterogeneous knowledge graph with an advanced Graph Attention Network framework (GATv2) to enhance phage-host interaction predictions. PHPGAT first constructs a knowledge graph by integrating phage–phage, host–host, and phage–host interactions, capturing complex relationships between biological entities. It then employs GATv2 to extract deep node features and learn dynamic interdependencies, generating context-aware embeddings. Finally, an inner product decoder computes the likelihood of interactions between phage and host pairs based on these embeddings.
@@
PHILM (Phage-Host Interaction Learning from Metagenomic profiles) is a approach developed to infer phage–host interactions (PHIs) directly from shotgun metagenomic data [490]. PHILM was trained to predict prokaryotic abundance profiles from phage abundance profiles, and then infers putative PHIs through a sensitivity analysis of the trained model. PHILM was evaluated in synthetic datasets using ecological models with known ground truth and curated paired phage–prokaryotic abundance profiles from real metagenomes. Across all evaluations, PHILM consistently outperformed traditional co-abundance methods in recovering true interactions. When applied to 7,016 healthy human gut metagenomes, PHILM identified 90% more genus-level PHIs than the assembly-based approach and revealed ecologically meaningful phage–host modules in the healthy gut. Moreover, the latent representations learned by PHILM captured microbial succession patterns in longitudinal data and served as powerful features for disease classification, outperforming taxonomic abundance-based features across multiple cohorts.
@@
Very recently, AI techniques have been applied to discover novel phage lysins [222, 223]. DeepLysin is a unified software package to employ AI for mining the vast genome reservoirs for novel antibacterial phage lysins [222]. DeepLysin consists of two modules: the lysin mining module and the antibacterial activity prediction module. The input of the lysin mining module is assembled contigs. This module consists of four steps: coding sequences (CDS) annotation, redundancy removal, domain alignment, and transmembrane protein removal. This module utilizes traditional blastP/protein sequence alignment-based methods to identify putative lysins. The second module estimates the antibacterial activity of the putative lysins identified by the first module. This module utilizes multiple AI techniques, such as Word2vec and an ensemble classifier that integrates five common classifiers to differentiate diverse and complex protein features. It ultimately applies Logistic Regression as a non-linear activation function to produce final activity predictions as scores ranging from 0 to 1, with higher scores indicating increased antibacterial activity. One limitation of DeepLysin is that four types of manually selected features (i.e., composition-based feature, binary profile-based feature, position-based feature, physiochemical based feature) need to be provided to the classifier. The feature selection procedure apparently heavily relies on domain knowledge.
@@ 4.1 Tradeoff between interpretability and complexity
Machine learning models, especially deep learning models, often suffer from high complexity and low interpretability, hindering their application in clinical decision-making. Those models are often referred to as black-box models, whose internal workings are hidden or not easily interpretable. In addition, deep learning models typically have more than thousands of neural weights whose training requires large sample sizes and high computational resources. We anticipate that those deep learning models can reach better performance than traditional machine learning models as long as the sample size is enough. However, in most clinic-related studies, traditional models (e.g., Random Forest) are still widely used due to their ease of implementation, smaller sample size requirement, and better interpretability.
@@
To address the interpretability issue, two different approaches can be employed. One approach is to employ methods such as SHAP (SHapley Additive exPlanations) [494], LIME (Local Interpretable Model-agnostic Explanations) [495] to enhance the interpretability of black-box machine learning models. SHAP is a game-theoretic method used to explain the output of any machine learning model. It links optimal credit allocation to local explanations by leveraging Shapley values from game theory and their related extensions. LIME is a technique that approximates any black box machine learning model with a local, interpretable model to explain each individual prediction. By applying SHAP and LIME, we can gain insights into complex deep learning models, identify biases, and improve transparency, crucial for applications in microbiome research.
@@
The other approach is to employ “white-box” models, (whose internal logic, parameters, and decision process are fully transparent and human-readable). For instance, ReduNet [496] is a white-box deep network based on the principle of maximizing rate reduction. The authors argued that, at least in classification tasks, a key objective for a deep network is to learn a low-dimensional, linearly discriminative representation of the data. The effectiveness of this representation can be assessed by a principled measure from (lossy) data compression, i.e., rate reduction. Appropriately structured deep networks can then be naturally interpreted as optimization schemes designed to maximize this measure. The resulting multi-layer deep network shares key characteristics with modern deep learning architectures, but each component of ReduNet has a well-defined optimization, statistical, and geometric interpretation. Applying ReduNet to microbiome data would be an interesting attempt. Unlike ReduNet, MDITRE is a supervised deep learning method specifically designed for microbiome research. It takes a phylogenetic tree, microbiome time-series data, and host status labels to learn human-interpretable rules for predicting host status [497]. The model consists of five hidden layers that can be directly interpreted in terms of if-then rule statements. The first layer focuses on phylogenetic relationships by selecting taxa relevant to predicting host status. The second layer focuses on time by identifying relevant time windows for prediction. The following layers determine whether the data from selected taxa and time windows exceed specific learned thresholds, and subsequently combine these conditions to generate the final rules for prediction.
@@ 4.3 Benchmarking evaluations
As we mentioned in previous sections several times, benchmarking evaluations are typically lacking in microbiology and microbiome research. Currently, there is no standardized pipeline for benchmarking machine learning or deep learning methods in microbiology and microbiome research. To ensure reproducibility across studies, it’s critical to standardize data preprocessing, which includes consistent methods for data collection, bioinformatics pipelines, and the profiling of microbiome taxonomies. Additionally, if feature dimension reduction is needed, it must be unbiased, using standardized methods for feature selection or reduction that apply uniformly across studies. Importantly, feature engineering should only be applied to training data and later evaluated on test data to avoid data leakage or overfitting. Furthermore, the creation of publicly available, well-annotated benchmarking datasets (analogous to MNIST or ImageNet in computer science) would provide the microbiome research community with reliable tools to assess and compare different machine learning models. Such datasets would accelerate progress and provide a framework for objective evaluation of new computational methods. Some attempts have been made in this regard. For example, MicrobiomeHD is a standardized database that compiles human gut microbiome studies related to health and disease [505]. It contains publicly available 16S data from published case-control studies, along with associated patient metadata. The raw sequencing data for each study were obtained and processed using a standardized pipeline. The curatedMetagenomicData package is another excellent example of benchmark microbiome datasets. It offers uniformly processed human microbiome data, including bacterial, fungal, archaeal, and viral taxonomic abundances, as well as quantitative metabolic functional profiles and standardized participant metadata [506]. This comprehensive, curated collection of metagenomic data is well-documented and easily accessible, making it suitable for benchmarking machine learning methods.
@@
Establishing benchmark datasets is critical for advancing AI application in microbiology and microbiome research. Such datasets can serve as standardized benchmarks that enable consistent, unbiased comparisons of algorithms, promote the development of robust predictive models, and foster reproducibility and transparency by allowing the research community to evaluate AI methods on a level playing field. Similar to the successful DREAM challenges in genomics, a community-driven effort to create public benchmarking datasets will foster collaboration, accelerate discovery, and establish best practices for AI approaches in microbiology and microbiome research. Collaborative input is vital for making this a reality.
4.4 Recent Breakthroughs
The rapid acceleration of AI development is widely recognized and driven by several converging factors—exponential increases in computing power, major global investments, and fundamental breakthroughs in algorithmic architectures. This unprecedented progress is reshaping scientific disciplines, and microbiology is no exception. We believe that microbiology and microbiome research stand to benefit profoundly from the expanding capabilities of modern AI systems. Below, we highlight two examples—AlphaFold 3 and Evo—that exemplify how generative and foundation models are beginning to transform our understanding of microbial life from sequence to ecosystem scale.
AlphaFold 3 [507] builds on the groundbreaking achievements of AlphaFold 2—recognized by the 2024 Nobel Prize for its impact on protein structure prediction—by extending its reach beyond individual proteins to entire biomolecular assemblies. AlphaFold 3 predicts 3D structures and interactions among proteins, DNA, RNA, small molecules, ions, and modified residues, effectively modeling “all of life’s molecules” at atomic resolution. It employs a streamlined generative pipeline: the Pairformer module encodes pairwise relationships between atoms across biomolecules using efficient attention mechanisms; these representations are then refined by a diffusion model that transforms noisy atomic coordinates into physically plausible 3D structures. A neural network backbone with triangular attention integrates evolutionary and geometric constraints, enabling end-to-end modeling of complex molecular interactions with remarkable speed and accuracy.
For microbiome science, AlphaFold 3 is poised to revolutionize our ability to model microbial communities at the molecular level. By predicting not only protein structures but also networks of protein–protein, protein–metabolite, enzyme–substrate, and host–microbe interactions, it provides a framework for translating metagenomic sequences into testable, mechanistic hypotheses. This capability may enable: (1) decoding functional pathways in unculturable bacteria, revealing novel enzymes for bioremediation or antibiotic biosynthesis; (2) mapping microbial drug targets at atomic precision, accelerating discovery of probiotics and phage therapies; (3) simulating microbe–host crosstalk by predicting how microbial ligands modulate immune receptors (e.g., TLRs, GPCRs); and (4) engineering synthetic microbial consortia based on protein-level compatibility and metabolic exchange. By bridging sequence data and structural function, AlphaFold 3 has the potential to shift microbiome research from correlative to causal, molecular-level understanding—unlocking precision manipulation of microbial ecosystems across health, agriculture, and environmental applications.
Evo [63] represents a complementary advance-a genomic foundation model trained on massive genomic corpora, including 2.7 million prokaryotic and phage genomes ($\approx 300$ billion nucleotide tokens) and later expanded to 9.3 trillion DNA base pairs from a highly curated genome atlas spanning all domains of life in Evo 2 [65]. It introduces a novel StripedHyena architecture that combines convolutional operators with rotary positional embeddings, allowing efficient long-context processing (up to one million nucleotides) with near-linear compute scaling. Trained via next-token prediction at single-nucleotide resolution, Evo learns the unified flow of biological information from DNA to RNA to protein—capturing motifs, regulatory grammar, and higher-order genome organization without task-specific fine-tuning.
In microbiology, Evo opens new avenues for predictive and generative genomics. It enables zero-shot identification of functional elements (e.g., promoters, splice sites, regulatory motifs), prediction of mutation effects, and de novo design of biological sequences. Applied to microbial systems, Evo could accelerate the design of synthetic microbes, precision-engineered phage cocktails, and enzymes for antibiotic resistance mitigation. For microbiome research, Evo’s multiscale modeling facilitates causal inference in community dynamics—linking genetic variation in unculturable gut bacteria to metabolic outputs, immune modulation (e.g., the interaction between short-chain fatty acids and toll-like receptors), or interspecies communication such as quorum sensing. When integrated with CRISPR-based editing and DNA synthesis platforms, Evo may ultimately enable programmable microbiome therapeutics and synthetic ecosystems for human and planetary health. As these capabilities expand, they also underscore the need for thoughtful consideration of biosafety and ethical governance in the era of AI-driven biological design.